Circular No. 7193

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V382 VELORUM M. Della Valle, European Southern Observatory (ESO) and Padova

University; L. Pasquini, ESO; and R. Williams, Max-Plank-Institut für Astrophysik, Garching, and Space Telescope Science Institute, report: “Preliminary analysis of spectra of N Vel 99 collected from May 28 to date, as part of a target-of-opportunity campaign at ESO with the 1.5-m telescope (+ Fiber-fed Extended Range Optical Spectrograph; range 390–900 nm, resolution 48 000), shows the continuum of the nova dominated by emission lines of the Balmer series with associated weak P-Cyg profiles, O I (multiplets 1, 4), Fe II (multiplets 27, 28, 38, 42, 48, 49, 74), N II, and Na I lines (D2 and D1 interstellar lines having EW = 0.039 and 0.037 nm). The [O III] forbidden-line identification reported by Hidayat et al. (IAUC 7188) at 500.7 nm is more likely to be the Fe II line (mult. 42) at 501.8 nm. The expansion velocities derived from the minima of the Balmer line P-Cyg profiles range between 2300 and 2400 km/s, whereas the HWZI measured for Hα yields 3600 km/s. The strong presence of Fe II lines, characterized by ‘flat top’ profiles and high expansion velocities, suggest this nova to have formed the lines in the discrete shell and thus to belong to the class of ‘broad’ Fe II novae (Fe II-b; see Williams 1992, A.J. 104, 725). The analysis of the preliminary lightcurve finds t_{2} ∼ 6 days and t_{3} ∼ 10 days. By combining these figures with the maximum magnitude vs. rate-of-decline relationship for classical novae (Della Valle and Livio 1995, Ap.J. 452, 704), we find an absolute magnitude at maximum of M_{V} ∼ −8.7 ± 0.2 (1σ) (cf. IAUC 7185).”

Further selected visual magnitude estimates: June 1.901 UT, 5.3 (T. A. Napoleao, São Paulo, Brazil); 3.902, 5.6 (W. C. de Souza, Sao Paulo, Brazil); 4.432, 5.7 (A. Pearce, Nedlands, W. Australia); 5.883, 6.1 (J. G. de S. Aguiar, Joaquim Egidio, Brazil); 6.436, 5.8 (Pearce); 7.439, 6.0 (Pearce).

SUPERNOVAE R. Szabo reports CCD R magnitude estimates of various supernovae

using the 1-m telescope at Piszkéstető, Hungary (based on USNO-A1.0 R magnitudes): SN 1999bg in IC 758, May 25.999 UT, 15.6; SN 1999bh in NGC 3435, May 25.936, 18.6; SN 1999br in NGC 4900, May 26.816, 16.8; SN 1999bt, May 26.027, 19.0; SN 1999bu in NGC 3786, May 25.965, 18.7; SN 1999bv in MCG +10-25-14, May 26.036, 18.0; SN 1999bw in NGC 3198, May 25.821, 18.5; SN 1999bx in NGC 6745, May 26.076, 19.0; SN 1999cb in Mkn 881, May 25.068, 17.7; SN 1999cc in NGC 6038, May 25.058, 16.7.

1999 June 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green